OpenStudy (anonymous):

Given the equation y − 4 = (x + 8) in point-slope form, identify the equation of the same line in slope-intercept form

OpenStudy (ghazi):

y-4=x+8 x-y=-12\[\frac{ x }{ -12 }+\frac{ y }{ -12 }=1\]